OpenStudy (anonymous):

Fraction Subtract and Divide 7/10-3/5 would it be 4/5 or 2/5

OpenStudy (kfujioka):

(7/10) - (3/5) = (7/10) - (6/10) = (7-6)/10 = 1/10